OpenStudy (anonymous):

What is the height of the triangle if the area is thirty-five square meters?

OpenStudy (anonymous):

5 meters 6 meters twenty meters twenty-one meters

OpenStudy (anonymous):

@lacrosseplayer22 @Venny

OpenStudy (anonymous):

5 meters

OpenStudy (cwrw238):

area = 0.5*base*height

OpenStudy (cwrw238):

38 = 0.5 * 14 * h

OpenStudy (cwrw238):

sorry - thats 35 not 38

OpenStudy (anonymous):

Area of a triangle = \({1\over2}\times{\rm base}\times{\rm height}\) \[35={1\over2}\times14\times h\]

OpenStudy (anonymous):

oh my bad i thought it was 7 not 7.5

OpenStudy (anonymous):

area=0.5*base*height. so 35 divided by 5 should get u the answer. h=5 becuz 35=1/2*14*h 35=7*h 35/7=h h=5 yw:) Army boy ur confusing me

OpenStudy (anonymous):

not to sorry simple error

OpenStudy (anonymous):

i got 5 but according to @cwrw238 im wrong

OpenStudy (anonymous):

A is the answer

OpenStudy (anonymous):

idk why he said that

OpenStudy (anonymous):

i got 5 but he said it comes out to 38

OpenStudy (anonymous):

@lacrosseplayer22 correct the answer is 5 meters

OpenStudy (anonymous):

twin i said that first look at top

OpenStudy (anonymous):

but @cwrw238 said i was wrong

OpenStudy (anonymous):

@Ail.A do you understand the steps now?

OpenStudy (anonymous):

@ArmyBoy97 but you didn't give step by step explanation.